Tohban Report 2015-07-29

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 22 Jul 2015
End Date: 29 Jul 2015
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: The one and only Juan Carlos Martinez Oliveros
List all reports


Solar Activity

Conditions are still very low activity, with only a handful of flares, but we did have some C class ones. It's expected that activity will stay quiet.

How many GOES flares occurred?

 Flares above B, C, M, X class were     13     4     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     11     1     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     0     0     0

There were RHESSI flares/GOES flares 42 / 17 over the time range 22-Jul-15 29-Jul-15

Memory Management

The memory is slowly emptying, now at ~6%. There was a Wallops hiccup that resulted in no Wallops passes on Saturday, so it's a great sign that we're getting down to a low SSR fill now. SSR max was 43%.

Spacecraft Status

Data Gaps

There were some few-hour gaps on July 23, maybe due to replays. This data is missing in the Browser. It's up to the tohban (outgoing, not incoming) to check the data archives and verify whether this data is actually present. If it's not there, she'll send out an alert.

Detector issues

Some thresholds were changed, though a few of them were a mistake on the tohban's part.

Detector 8 rear fast threshold was raised all the way to the max, in order to try and recover that segment's performance. This was unsuccessful, and there continue to be no useful data returned from detector 8 rear.

Detector 4 front fast threshold was raised to free up livetime; this was successful, with livetime now around 90%.

Detectors 1, 6, and 7 rear fast threshold was changed due to low livetime, but it turned out this change was unnecessary. The tohban had not noticed that the previous tohban's changes had not yet "come through" in observable results due to the time lag in getting the data with the memory not emptying. The next tohban should decide whether to reverse these changes.

The command log:

2015-205-17:19:08 /IDPLOAD VALUE=0x70 ; was 0x60
2015-205-17:19:45 /IDPLOAD VALUE=0xB0 ; was 0xA0
2015-205-17:20:08 /IDPLOAD VALUE=0x70 ; was 0x60
2015-205-17:21:50 /IDPLOAD VALUE=0xFF ; was 0xD0

2015-205-17:22:30 /IDPLOAD VALUE=0x90  ; was 0x80

Other notes


Both D1 and D7 livetimes had continued to degrade since we last changed the attenuator deinsertion settings. Livetime values are now flipping between 97.25% and 97.65% and only rarely jumping a level up from that. Attenuator deinsertion levels at the beginning of this tohban week were 98.04% for the thin and 98.43% for the thick, so if either attenuator were to come in it would not come out until commanded to. Therefore, a change in deinsertion threshold was requested. The A3 out threshold was change from 1.57% to 2.75% deadtime and the A1 out threshold was changed from 1.96 to 2.75% deadtime. The changes were made today, July 24, at 2015-204-17:44:08 UT. The last attenuator motion (A1 to A0) was July 9, 2015.

There were two thin attenuator motions over the weekend, both nominal.


Spacecraft Management

Decimation Normal/Vigorous
Night time data (fronts) +- 4 minutes
Night time data (rears) +- 4 minutes
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools